### 16 Apr 2026 v2

# BlasBench: An Open Benchmark for Irish Speech Recognition github.com/jyoutir/blasbench

Abstract
Existing multilingual benchmarks include Irish among dozens of languages but apply no Irish-aware text normalisation, leaving reliable and reproducible ASR comparison impossible. We introduce BlasBench, an open evaluation harness that provides a standalone Irish-aware normaliser preserving fadas, lenition, and eclipsis; a reproducible scoring harness and per-utterance predictions released for all evaluated runs. We pilot this by benchmarking 12 systems across four architecture families on Common Voice ga-IE and FLEURS ga-IE. All Whisper variants exceed 100% WER through insertion-driven hallucination. Microsoft Azure reaches 22.2% WER on Common Voice and 57.5% on FLEURS; the best open model, Omnilingual ASR 7B, reaches 30.65% and 39.09% respectively. Models fine-tuned on Common Voice degrade 33–43 points moving to FLEURS, while massively multilingual models degrade only 7–10—a generalisation gap that single-dataset evaluation misses.
BlasBench: An Open Benchmark for Irish Speech Recognition
 github.com/jyoutir/blasbench
Jyoutir Rajthanks: Corresponding author. Independent Researcher jyoutirraj@gmail.com          John Conway Independent Researcher john@blasapp.com
1 Introduction
State-of-the-art multilingual systems fail on out-of-distribution audio in Irish. Microsoft Azure reaches 22.2% WER on read speech (Common Voice ga-IE) and 57.5% WER on the more naturalistic FLEURS ga-IE. Every Whisper variant we evaluate (Radford et al., 2023) produces more tokens than the reference contains, exceeding 100% WER on both datasets (Tables 12). Irish presents a genuine difficulty: initial consonant mutations, eclipsis, and a spelling system that encodes grammatical contrasts mean that language-agnostic normalisation silently discards linguistically meaningful variation.
The problem is not that nobody has worked on Irish ASR. The ABAIR group at Trinity College Dublin has produced a sustained research programme culminating in their Fotheidil system at 10.9% WER (Lonergan et al., 2025). Qian et al. (2024) fine-tune Whisper on Irish. Community wav2vec2 fine-tunes exist on HuggingFace (Faste, 2022). The IWSLT shared task has produced Irish–English speech translation systems (Ahmad et al., 2024; Moslem, 2024). Multilingual benchmarks—FLEURS (Conneau et al., 2022a), MMS (Pratap et al., 2023), ML-SUPERB (Shi et al., 2023)—include Irish among dozens or hundreds of languages. These efforts span architecture families and data regimes, yet they have not been placed on a common footing.
The gap is methodological rather than empirical. ABAIR’s best result is on private MíleGlór data; their published Common Voice and FLEURS-R figures cover only their own systems. Community models self-report on different Common Voice versions with different normalisers. Multilingual benchmarks include Irish without Irish-specific normalisation or targeted analysis. [^1] ML-SUPERB and XTREME-S (Conneau et al., 2022b) include Irish data points but do not compare end-user ASR systems on Irish specifically. We did not identify a prior open, Irish-specific comparison of end-user ASR systems across architecture families under a shared Irish-aware evaluation protocol.
We present BlasBench, [^2] Harness: github.com/jyoutir/blasbench. Per-utterance predictions and aggregate results for every run in this paper are released as v0.1.0. an open benchmark and evaluation harness for Irish ASR. The harness applies Irish-aware text normalisation before scoring, so that orthographic variants reflecting mutation and dialect do not inflate error rates artificially. A 12-model comparison across commercial, large-scale multilingual, and fine-tuned systems constitutes the first use of the benchmark; reproducible, language-aware scoring is also a prerequisite for any automated research pipeline that requires a stable signal. Dedicated low-resource benchmarks for Urdu (Arif et al., 2025) and Scottish Gaelic (Klejch et al., 2025) have made the same case for their respective languages; BlasBench brings it to Irish.
2 The BlasBench Benchmark
2.1 Task and scope
BlasBench evaluates Irish speech-to-text systems. Given Irish audio, a system produces Irish text; the benchmark scores the output against a reference transcription using WER and CER, computed with Irish-aware normalisation.
2.2 Datasets
Two public datasets, chosen to measure both in-distribution and out-of-distribution performance:
  • Common Voice 25.0 ga-IE (Ardila et al., 2020): 874 community-recorded read-speech test utterances in the released BlasBench evaluation slice; the utterance count is recorded in the released run metadata. Most wav2vec2 fine-tunes we evaluate were trained on earlier CV versions.
  • FLEURS ga-IE (Conneau et al., 2022a): 842 test utterances in the released BlasBench evaluation slice; the utterance count is recorded in the released run metadata. The speech is professionally recorded native-speaker read speech. Based on public model cards, no model in our benchmark reports FLEURS in its training mix.
The released prediction files and run metadata make each evaluation slice auditable. Using two datasets with different recording conditions is central to the benchmark design: it exposes generalisation failures invisible to single-dataset evaluation (§5.2).
2.3 Evaluation pipeline
Audio input (16kHz)Model wrapper (any architecture)Irish-aware normaliserGlobal WER / CERBootstrap 95% CIraw text
Figure 1: BlasBench evaluation pipeline. The Irish normaliser (highlighted) is the language-specific component.
The pipeline (Figure 1) proceeds in four stages:
  1. 1.
    Inference. Audio is passed through a model wrapper that returns raw text. The wrapper API is standardised: any system that accepts 16kHz audio and returns a string can be evaluated.
  2. 2.
    Normalisation. Both reference and hypothesis are normalised using an Irish-aware pipeline (§2.4).
  3. 3.
    Scoring. WER and CER are computed via global aggregation: total substitutions, insertions, and deletions across all utterances divided by total reference units. This corpus-level approach avoids the bias of per-sentence averaging and matches the convention used in low-resource benchmarking work (Arif et al., 2025).
  4. 4.
    Confidence intervals. Bootstrap 95% CIs (Bisani and Ney, 2004) are computed with 1,000 resamples at a fixed seed (42), resampling utterance-level error counts and recomputing the global aggregate. CIs for every reported point estimate are released alongside the predictions in each run’s results.json; tables show point estimates for compactness.
2.4 Irish-aware normalisation
Standard ASR normalisers lowercase text, strip punctuation, and collapse whitespace. Recent work on multilingual ASR evaluation shows that language-agnostic normalisation can materially distort reported error rates (Manohar et al., 2024). For Irish, it destroys linguistically significant information.
Irish uses acute accents (fadas) on five vowels: á, é, í, ó, ú. These are phonemically contrastive: fear “man” vs féar “grass.” Standard NFKD normalisation decomposes accented characters and may strip the combining accent. Our normaliser applies NFC first, guaranteeing fadas survive.
Irish also marks grammar through initial mutations. Lenition: bean “woman” $\to$ bhean “his wife.” Eclipsis: bean $\to$ mbean “of women.” Our normaliser preserves these.
ABAIR’s pipeline includes text-processing components used inside Fotheidil (Lonergan et al., 2025) but no Irish ASR normaliser has been released standalone. To our knowledge, BlasBench ships the first standalone open-source Irish ASR evaluation normaliser released specifically for Irish ASR scoring; unreleased internal tools may exist.
2.5 Released artifacts
For each (model, dataset) run, BlasBench produces:
  • Per-utterance predictions with sample ID, reference, hypothesis, WER, CER
  • Aggregate results with S/I/D breakdown and bootstrap CIs
  • Run metadata recording dataset name/split/count, model identity, evaluation configuration, and software versions
The evaluation harness and released predictions are available at the repository linked under the title. Future systems can be compared against these predictions without re-running prior models.
2.6 Reproducibility and extensibility
Adding a new model requires implementing a single function: given 16kHz audio, return a string. The harness handles normalisation, scoring, CI computation, and artifact generation.
3 Systems Evaluated
We evaluate 12 systems spanning four architecture families.
Whisper (Radford et al., 2023): Encoder-decoder transformer. Irish is absent from Whisper’s language list across v1, v2, and v3,333ga is absent from the LANGUAGES dictionary in whisper/tokenizer.py for v1/v2/v3, and the v3 model card provides no per-language training-data disclosure that would indicate Irish coverage. so results here are effectively zero-shot. We test medium (769M), large-v2 (1.5B), large-v3 (1.5B), and large-v3-turbo (809M).
wav2vec2 CTC: Self-supervised XLS-R or XLSR-53 encoders with CTC heads, fine-tuned on Irish Common Voice by community contributors (315M–1B).
Meta multilingual: MMS-1B-All (Pratap et al., 2023) (1,107 languages); Omnilingual ASR (omniASR) LLM (Keren et al., 2025), wav2vec2 encoders with LLM-inspired Transformer decoders covering 1,600+ languages, released in 300M, 1B, 3B, and 7B sizes; we test the 300M and 7B endpoints.
Commercial: Microsoft Azure Speech Services (ga-IE locale).
Open-weights models ran on a single NVIDIA H100 80GB SXM via RunPod ($\sim$6 GPU hours, $\sim$\$25). Azure ran via API ($\sim$\$5 at April 2026 pricing). Whisper, wav2vec2, MMS, and Azure runs all used the released harness; the omniASR runs used a separate fairseq2 driver at the time of writing, with a harness adapter pending.
4 Results
Tables 1 and 2 present WER, CER, and per-error-type breakdown on both datasets.
Table 1: Common Voice ga-IE (874 utterances). ABAIR is self-reported (Lonergan et al., 2025) using a different normaliser. Whisper results are zero-shot. Type column: open = weights and training data public; open-w = weights public, training data not fully disclosed; API = commercial cloud API; closed = weights unavailable.
[…]
ref ABAIR / Fotheidil closed TDNN-HMM 19.6*
1 azure / speech-ga-IE API proprietary 22.2 15.8 1.7 4.8 11.4
2 omniASR LLM 7B open-w w2v2+Trans 30.6 25.0 2.5 3.2 14.6
3 Aditya3107 / xls-r-1b open w2v2 CTC 32.4 26.4 1.7 4.3 12.8
4 omniASR LLM 300M open-w w2v2+Trans 37.6 29.3 3.4 4.9 19.2
5 kingabzpro / xls-r-1b open w2v2 CTC 45.8 38.2 3.6 4.0 18.9
6 jimregan / xlsr-53 open w2v2 CTC 48.9 40.5 4.2 4.2 20.3
7 cpierse / xlsr-53 open w2v2 CTC 49.4 41.5 3.9 4.0 21.0
8 mms-1b-all open w2v2 CTC 54.2 44.1 2.8 7.4 21.2
9 whisper-large-v2 open enc-dec 106.0 73.6 19.9 12.4 68.5
10 whisper-large-v3 open enc-dec 125.6 78.8 33.1 13.7 85.6
11 whisper-medium open enc-dec 129.3 76.2 40.6 12.5 91.8
12 whisper-large-v3-turbo open enc-dec 225.6 83.1 128.8 13.7 159.7
Table 2: FLEURS ga-IE (842 utterances). ABAIR’s 44.5% is on FLEURS-R (Ma et al., 2024), not this test set.
[…]
ref ABAIR / Fotheidil closed TDNN-HMM 44.5*
1 omniASR LLM 7B open-w w2v2+Trans 39.1 32.2 3.4 3.5 18.6
2 omniASR LLM 300M open-w w2v2+Trans 47.7 38.4 4.9 4.4 24.1
3 azure / speech-ga-IE API proprietary 57.5 21.5 3.5 32.5 43.8
4 mms-1b-all open w2v2 CTC 61.6 51.9 3.2 6.5 26.0
5 Aditya3107 / xls-r-1b open w2v2 CTC 75.8 62.1 6.2 7.5 36.0
6 kingabzpro / xls-r-1b open w2v2 CTC 78.5 64.7 9.2 4.6 38.2
7 jimregan / xlsr-53 open w2v2 CTC 83.0 68.2 9.8 5.0 40.8
8 cpierse / xlsr-53 open w2v2 CTC 83.2 68.6 9.8 4.8 41.9
9 whisper-large-v2 open enc-dec 102.8 78.2 19.8 4.8 59.6
10 whisper-medium open enc-dec 134.1 86.5 43.1 4.4 87.3
11 whisper-large-v3 open enc-dec 217.8 89.8 123.7 4.3 156.0
12 whisper-large-v3-turbo open enc-dec 587.6 91.2 491.2 5.1 410.1
5 Benchmark Findings
5.1 The Whisper variants we evaluated fail catastrophically on Irish
All four Whisper variants exceed 100% WER on both datasets, with insertion rates of 20–491%: the decoder emits fluent English unrelated to the input (Appendix C). v3 is worse than v2 (125.6% vs 106.0% on CV), and turbo is worse again (587.6% on FLEURS). Qian et al. (2024) report 110.4% WER for whisper-large-v3 on FLEURS Irish, which is in the same regime as our results; we did not find prior documentation of the monotonic v2 $\to$ v3 $\to$ turbo regression on Irish.
5.2 Common Voice overestimates performance
Table 3: Cross-corpus generalisation gap. $\Delta$ = FLEURS - CV WER. Multilingual models generalise; CV-trained models do not.
model cv fleurs $\Delta$
$m$s-1b-all 54.2 61.6 +7.3
omniASR 7B 30.6 39.1 +8.4
omniASR 300M 37.6 47.7 +10.1
azure 22.2 57.5 +35.2
cpierse 49.4 83.2 +33.8
Aditya3107 32.4 75.8 +43.4
Models fine-tuned on Common Voice degrade 33–43 WER points on FLEURS (Table 3). Azure degrades 35 points. Models with massively multilingual pre-training degrade only 7–10 points. Prior work has evaluated across corpora (Lonergan et al., 2025), but the magnitude of this gap has not been quantified under a single harness. Common Voice WER alone is not a reliable proxy for deployment; BlasBench therefore requires evaluation on at least two datasets.
5.3 The open–closed gap
ABAIR’s Fotheidil reports its best results (10.9% WER on private MíleGlór, 19.6% on Common Voice, and 44.5% on FLEURS-R) using their full pipeline including RNNLM rescoring; without LM rescoring the M1 acoustic model alone reports 23.7% on Common Voice (Lonergan et al., 2025; Ma et al., 2024). [^4] ABAIR evaluates on FLEURS-R (audio restored from the original FLEURS via the Miipher model), not the original FLEURS we use, and with a different normaliser. Numbers are indicative, not directly comparable. omniASR 7B reaches 30.6% and 39.1% on original FLEURS under our normaliser. Because datasets and normalisers differ, we treat these numbers as indicative rather than a direct open-vs-closed measurement; they are consistent with a narrower gap than the headline 10.9%/30.6% suggests, but do not establish one. ABAIR’s architecture (TDNN-HMM, Kaldi-based) is from a different family than the strongest open models here (wav2vec2-based), and their reported gains coincide with substantial Irish-specific training data (398h labelled, 3,230h pseudo-labelled (Lonergan et al., 2025)) that exceeds what any open model was trained on for Irish. The primary binding constraint on Irish ASR is data, not architecture.
6 Related Work
Irish ASR.
ABAIR has published on hybrid systems (Lonergan et al., 2022a), dialect lexicons (Lonergan et al., 2022b), dialect-balanced training (Lonergan et al., 2023a), dialect identification (Lonergan et al., 2023b), multi-task learning (Lonergan et al., 2024), and Fotheidil (Lonergan et al., 2025). Qian et al. (2024) fine-tune Whisper for Irish but compare only Whisper variants. To our knowledge, no prior work compares multiple end-user Irish ASR systems from different architecture families under a single Irish-aware evaluation protocol.
Low-resource and Celtic-language benchmarks.
Arif et al. (2025) benchmark Whisper, MMS, and SeamlessM4T on Urdu with corpus-level WER aggregation; BlasBench adopts a similar aggregation strategy and adds bootstrap CIs and cross-corpus analysis. Klejch et al. (2025) build Scottish Gaelic ASR with Gaelic-specific normalisation, Jones (2022) fine-tune wav2vec2 for Welsh, and Bartley and Ragni (2025) demonstrate minimal-data ASR for Manx and Cornish from a spoken dictionary. On evaluation methodology, Manohar et al. (2024) show that multilingual ASR normalisers can distort error rates when they erase language-specific orthographic information. On statistical methodology, Gillick and Cox (1989) and Bisani and Ney (2004) remain the canonical references for significance testing of WER comparisons.
Multilingual foundation models and benchmarks.
Beyond MMS and Whisper, Zhang et al. (2023) (Google USM), Puvvada et al. (2024) (NVIDIA Canary), and Peng et al. (2024) (OWSM-CTC) all target broad multilingual speech coverage, but none of these papers establishes end-user Irish ASR performance. Multilingual benchmarks FLEURS (Conneau et al., 2022a), ML-SUPERB (Shi et al., 2023), and XTREME-S (Conneau et al., 2022b) include Irish among many languages but apply no Irish-specific normalisation; the Open ASR Leaderboard (Srivastav et al., 2025) does not include Irish at all (its multilingual track currently covers German, French, Italian, Spanish, and Portuguese). Barry et al. (2022) release gaBERT for Irish NLP but not for ASR evaluation.
7 Conclusion
BlasBench is, to our knowledge, the first open Irish-specific benchmark that compares end-user ASR systems under a shared Irish-aware protocol. Across 12 systems on two datasets, wav2vec2 models are the only viable family we tested, the Whisper variants we evaluate fail catastrophically, and the gap between Common Voice and FLEURS scores exposes a generalisation failure that single-dataset evaluation hides. The deeper bottleneck is data: labelled Irish audio remains the binding constraint on progress. Because scoring is deterministic and reproducible, BlasBench is suitable as an evaluation layer within automated research pipelines (Karpathy, 2026), where reliable comparison across many candidate models depends on a stable scoring function. The harness, normaliser, per-utterance predictions, and run metadata are publicly released.
Limitations
We evaluate on fixed public BlasBench slices for Common Voice ga-IE (874 utterances) and FLEURS ga-IE (842 utterances); both consist of read speech, and no public conversational Irish ASR benchmark exists. These sample sizes limit the precision of pairwise comparisons. Confidence intervals are reported in the accompanying JSON; main tables show point estimates for compactness, as noted in the evaluation pipeline.
Neither dataset spans all three dialects equally, and neither provides dialect labels, reflecting a gap in available Irish data rather than a design choice. Dialect-stratified evaluation is not possible with current public resources.
The 12 systems evaluated cover the main open, community, and commercial categories; selection was constrained by licence, cost, and available credentials. OmniASR was run outside the shared evaluation harness because a fairseq2 driver dependency was unresolved at the time of writing, with a harness adapter pending; its numbers are therefore not directly comparable at the implementation level. Azure results depend on the API version and model snapshot available at time of evaluation (April 2026); future versions may differ.
ABAIR uses FLEURS-R rather than original FLEURS as its test set, and its normaliser is not publicly available; the open–closed comparison is therefore indicative rather than controlled. The WER impact of our normalisation choices has not been quantified.
The community wav2vec2 models were fine-tuned on earlier Common Voice releases (CV 6.1–8.0) and evaluated on CV 25.0. Some of the cross-corpus gap likely reflects distribution shift alongside genuine generalisation behaviour, with the multilingual-pretrained models in Table 3 providing a partial control.
Acknowledgements
Portions of this work were assisted by Claude Opus 4.6 (Anthropic, 2026).
Appendix A Error-Type Breakdown
WER alone conflates three failure modes that point at different fixes. Table 4 reports S, I, and D as percentages of total reference words, aggregated globally.
Table 4: Error-type breakdown. S, I, D reported as percentages of reference words, aggregated globally. Sorted by WER within each dataset.
model wer s i d
Common Voice
azure ga-IE 22.3 15.8 1.7 4.8
omniASR 7B 30.7 25.0 2.5 3.2
Aditya3107 xls-r-1b 32.4 26.4 1.7 4.3
omniASR 300M 37.6 29.3 3.4 4.9
kingabzpro xls-r-1b 45.8 38.2 3.6 4.0
jimregan xlsr-53 48.9 40.5 4.2 4.2
cpierse xlsr-53 49.4 41.5 3.9 4.0
$m$s-1b-all 54.3 44.1 2.8 7.4
whisper-large-v2 106.0 73.6 20.0 12.4
whisper-large-v3 125.6 78.8 33.1 13.7
whisper-medium 129.3 76.2 40.6 12.5
whisper-large-v3-turbo 225.6 83.1 128.8 13.7
FLEURS
omniASR 7B 39.1 32.2 3.4 3.5
omniASR 300M 47.7 38.4 4.9 4.4
azure ga-IE 57.5 21.5 3.5 32.5
$m$s-1b-all 61.6 51.9 3.2 6.5
Aditya3107 xls-r-1b 75.8 62.1 6.2 7.5
kingabzpro xls-r-1b 78.5 64.7 9.2 4.6
jimregan xlsr-53 83.0 68.2 9.8 5.0
cpierse xlsr-53 83.2 68.6 9.8 4.8
whisper-large-v2 102.8 78.2 19.8 4.8
whisper-medium 134.1 86.5 43.1 4.4
whisper-large-v3 217.8 89.8 123.7 4.3
whisper-large-v3-turbo 587.6 91.2 491.2 5.1
wav2vec2 errors are substitutions; Whisper errors are insertions. Every wav2vec2 model keeps I below 10%; every Whisper variant exceeds 20%, peaking at 491% for large-v3-turbo on FLEURS. That is not an acoustic failure (see Appendix C). Going CV $\to$ FLEURS, wav2vec2 S and I both roughly double while D stays flat; Whisper I grows another order of magnitude—longer audio gives the decoder more runway to hallucinate. Azure is the anomaly: S-dominated on CV, D-dominated on FLEURS (4.8 $\to$ 32.5), suggesting a VAD or endpointer tuned for short clips.
Appendix B What the Hard Utterances Look Like
Filtering to utterances where every non-Whisper model exceeds 50% WER isolates text-level difficulty. Two distinct tails emerge.
Common Voice hard cases are short (4–6 words, where one error is 15–25 WER points by construction) and Irish-internal: dialect forms (chuile, bréidín, comhluadar, airneáin), copula and conditional idioms (gréasaí ab ea é, ní chuirfinn thairis é), and Gaeltacht place names in their eclipsed form (i ngaillimh thiar, leitir mealláin). Hard because training data underrepresents them, not because the text is exotic.
FLEURS hard cases are code-switched news text: alphanumerics fused into single tokens (80211n, 24ghz, m16, 10001100), foreign proper nouns (wong kan seng, rolanda mendoza, aerosmith), English-origin acronyms pronounced as English (mdt, swapo), and bare in-text academic citations (larson agus lafasto 1989 lch 109). A number-and-acronym expansion rule would partially fix this tail.
Appendix C Representative Whisper Outputs
Four examples showing what Whisper’s insertions actually are. ref = Irish reference, w3 = whisper-large-v3, wv2 = Aditya3107/xls-r-1b (acoustically weak, does not hallucinate).
(a) Whisper emits Welshcv 545
ref dia dhaoibh tráthnóna
w3 diolch yn fawr iawn am wylior fideo
wv2 dia dhaoibh tráthnóna
(b) Repetition loopcv 216 (turbo)
ref tabhair cabhair don fhoireann
w3 to a coward to a coward to a coward …(333 tokens)
wv2 tabhair cabhair don fhoireann
(c) Unrelated English sign-offfl 609
ref phléasc buama amháin lasmuigh doifig an ardghobharnóra
w3 thank you for listening and have a good day
wv2 pléis buam amhain leasmúid duifigh an ard gabhrana
(d) Topic-adjacent Englishfl 98 (turbo)
ref ina dhiaidh sin bogadh chuig ospidéal addenbrooke i gcambridge é
w3 in the next day ill be back to edinburghs hospital in cambridge
wv2 ina dhíg sin bothar chuig ospadéal adan bhrog a ceamraid é
Case (d) is the telling one: Whisper catches cambridge, hears addenbrooke as edinburgh, and wraps them in fluent English syntax. The acoustic signal reaches the decoder; the decoder refuses to emit Irish. Irish is a rounding error in Whisper’s token budget (§5.2), so its strongest prior on any audio segment is a language it knows.